# Reproducing the first and second moment of empirical degree distributions

### v3

@@ Abstract
The study of probabilistic models for the analysis of complex networks represents a flourishing research field. Among the former, Exponential Random Graphs (ERGs) have gained increasing attention over the years. So far, only linear ERGs have been extensively employed to gain insight into the structural organisation of real-world complex networks. None, however, is capable of accounting for the variance of the empirical degree distribution. To this aim, non-linear ERGs must be considered. After showing that the usual mean-field approximation forces the degree-corrected version of the two-star model to degenerate, we define a fitness-induced variant of it. Such a ‘softened’ model is capable of reproducing the sample variance, while retaining the explanatory power of its linear counterpart, within a purely canonical framework.
@@ I Introduction
A class of models that has gained increasing attention over the years is that of Exponential Random Graphs (ERGs) [10, 16, 17, 11, 12, 18, 13, 14, 15]. ERGs belong to the category of canonical approaches, being induced by constraints that are soft, i.e. can be violated by individual configurations even if their ensemble average matches the enforced value exactly. A key advantage of canonical approaches is that the expected value of topological properties can be often expressed analytically in terms of the constraints, thereby avoiding the computational cost of generating randomised networks [19]. Microcanonical approaches, instead, artificially generate many randomised variants of the observed network, enforcing constraints that are hard, i.e. met exactly by each graph in the ensemble [20, 21, 22, 23, 24]; this strong requirement, however, comes at the price of non-ergodicity, high computational demand and poor generalisability [25, 26].
So far, only linear ERGs have been extensively employed to gain insight into the structural organisation of real-world systems: for simple graphs, the most important one is the Undirected Binary Configuration Model (UBCM), inducing an ensemble of configurations specified by the degree sequence; a useful fitness-induced variant of it, to be employed in presence of partial information is, instead, the density-corrected Gravity Model (dcGM), solely enforcing (a proxy of) the link density while relying on node-specific quantities identified with the node strengths $\{s_{i}\}_{i=1}^{N}$, defined as $s_{i}=\sum_{j(\neq i)}w_{ij}$, $\forall\>i$ and representing the total volume of interactions involving each node. This choice is motivated by the evidence that in economic and financial applications a node strength is observable from balance-sheets, transactional or input-output records, while the number of its counterparties (i.e. its degree) is typically not disclosed [27, 28, 29, 30, 31, 32]: treating strengths as exogenous fitnesses, thus, allows one to account for the heterogeneity of nodes without imposing any local constraint [19]: in formulas,
@@
The aim of this work is to investigate whether it is possible to define a minimal model capable of accurately reproducing both the first and second moment of empirical degree distributions, without resorting to microcanonical constraints. To this end, we consider the class of ERGs with non-linear constraints and introduce a fitness-induced variant of the two-star model that can be efficiently and reliably implemented. We test it on the transaction-level data constituting the overnight segment of the Electronic Market for Interbank Deposits (eMID), a screen-based market for unsecured deposits [8, 38, 39]. Although the raw records are directed and weighted (lender, borrower, notional amount), here we symmetrise and binarise exposures within each, considered time window, i.e. ‘daily’, ‘weekly’, ‘monthly’, ‘quarterly’ and ‘yearly’ (the description of each aggregation procedure is reported in Appendix A).
@@ II The simplest non-linear constraint
The simplest non-linear constraint is represented by the total number of two-stars [10, 40], i.e.
@@
Figure 1: Analysis of eMID during the 31st week (left panel) and the 42nd week (right panel) of the year 2004. Graphical representation of the agreement between the sample variance $\text{Var}[\bm{k}]$ (black, dashed, vertical line) and its expected value $\langle\text{Var}[\bm{k}]\rangle$ under the UBCM (red), the dcGM (yellow) and the fit2SM (brown): while the UBCM steadily overestimates it and the dcGM either overestimates or underestimates it, the fit2SM correctly reproduces it. Each ensemble distribution is well approximated by a Gaussian whose parameters match the corresponding average and standard deviation.
@@ III A microcanonical approach?
Since the result above suggests the microcanonical road as the only viable one, a natural choice would be that of considering the microcanonical version of the two-star model (m2SM), constraining both $L$ and $S$ exactly: since $\langle L\rangle=L$, $\text{Var}[L]=0$ and $\langle S\rangle=S$, one would obtain $\langle\text{Var}[\bm{k}]\rangle=\text{Var}[\bm{k}]$. This model is, however, homogeneous and, therefore, unable to reproduce the degree sequence.
@@
satisfying both (sets of) constraints microcanonically is equivalent to require that $\text{Var}[k_{i}]=0$, $\forall\>i$. In words, our results indicate that the microcanonical version of the Configuration Model (mCM) is the (simplest) one guaranteeing $\langle k_{i}\rangle=k_{i}$, $\forall\>i$ and $\langle S\rangle=S$. Moreover, as the relationship
@@
While the (microcanonical) requirements are clear, the way to realise them is such to a much lesser extent; moreover, the problems affecting the microcanonical approaches mentioned in the introductory paragraph let us opt for not taking such a road. Is a canonical way out viable?
@@ IV The mean-field approximation
the dc2SM is, however, unable to match both (sets of) constraints, as reproducing the empirical number of two-stars implies underestimating the degree of at least one node (see also Appendix C). The explanation of such a behaviour lies in a simple observation: as the mean-field approximation scheme requires that
@@
letting eq. III vanish requires the generic probability coefficient of a canonical model to degenerate, becoming either 0 or 1: in other words, constraining both the degrees and the number of two-stars within the mean-field approximation scheme forces the model to become deterministic, individuating the observed configuration [^1] The asterisk indicates (the value of the quantities measured on) the observed configuration $\mathbf{A}^{*}$. $\mathbf{A}^{*}$ as the only admissible one.
@@ V A canonical way out
Let us, now, ask ourselves if an alternative, canonical road exists. To this aim, let us, first, notice that $\text{Var}[L]$ is divided by $N^{2}$, in eq. II: such an addendum may, thus, be expected not to play a relevant role. As a consequence, we are allowed to consider the expression
@@
Figure 3: Left panel: distributions of the spectral radius induced by $M=10^{3}$ configurations sampled from the Chung-Lu model defined by the topology of the eMID snapshot corresponding to the day 20102010-0707-1919, according to the UBCM (red), the dcGM (yellow) and the fit2SM (brown). The three models were solved for each of the $M=10^{3}$, generated configurations and the corresponding ensembles explicitly sampled in order to obtain an estimation of $\pi_{1}$: what we find is that $\lambda_{1}^{\text{dcGM}}<\pi_{1}\lesssim\lambda_{1}^{\text{fit2SM}}<\lambda_{1}^{\text{UBCM}}$. In words, while the UBCM overestimates the spectral radius of the generative model (black, dashed, vertical line) and the dcGM underestimates it, the fit2SM correctly reproduces it. Each ensemble distribution is well approximated by a Gaussian whose parameters match the corresponding average and standard deviation. Right panel: violin plots (summarising the distributions) of the relative error in reproducing the empirical spectral radius $\lambda_{1}^{\text{obs}}$, across snapshots and temporal aggregations. The fit2SM (brown) yields smaller errors i) than the dcGM (yellow) at all time-scales; ii) than the UBCM (red) at the daily and weekly time-scales.
Constructing a fitness-induced variant of the 2SM (fit2SM) is, nevertheless, feasible: to this aim, it is enough to replace $x_{i}$ with $\sqrt{z}s_{i}$ [^3] To ease numerical manipulations, we have rescaled the strengths, dividing them by their arithmetic mean. and $k_{i}$ with its expectation under the same model, in eq. 10. These substitutions lead to the generic probability coefficient reading
@@
where $\langle k_{i}\rangle_{\text{fit2SM}}\equiv\kappa_{i}$. The two, global parameters $z$ and $y$ can be estimated by solving the (system constituted by the) two, non-linear, coupled equations reading
@@
Figure 4: For each snapshot and aggregation level, we show the violin plots (summarising the distributions) of the relative error $\delta\text{BIC}_{model}$, defined in eq. 27. Upon remembering that a null value of such an index identifies the best-performing model on the considered snapshot, let us notice that the score induced by the fit2SM (brown) concentrates near zero - hence such a model dominates over the competing ones - at finer aggregations, while the UBCM (red) is the winner at coarser aggregations.
@@ VI Testing the performance
VI.1 Reproducing a network local properties
Since the degree sequence is a standard diagnostic in network reconstruction exercises, we start our analysis by testing the accuracy of the fit2SM in reproducing it. We do so by means of the quantile-quantile (QQ) plot: given the empirical sequence $\{k_{i}\}_{i=1}^{N}$ and a model-based, expected one $\{\langle k_{i}\rangle\}_{i=1}^{N}$, we sort both in increasing order and scatter the corresponding pairs of values [^4] Recall that the quantile $q$ of a set of (ordered) values indicates the percentage $q$ of smallest values. Given the empirical degree sequence and the model-based, expected one, we, thus, compute their quantile functions by sorting both in increasing order and scattering the $r$-th order statistics $\langle k\rangle_{r}$ versus $k_{r}$, $r=1\dots\>N$.; while perfect agreement would place all points on the identity line, mis-estimations of the values would be revealed by systematic deviations.
Let us, now, select five snapshots by drawing one calendar year at random and, conditionally on it, drawing one quarter, one month, one ISO week and one trading day at random as well. Figure 2 reports the resulting QQ plots for year 2001, quarter Q3, month April, ISO week 2 and day 20012001-1111-2121; the visual inspection is accompanied by the accuracy indicators named average relative error, defined as $\text{ARE}=N^{-1}\sum_{i}|k_{i}-\langle k_{i}\rangle|/k_{i}$, and maximum relative error, defined as $\text{MRE}=\max_{i}\{|k_{i}-\langle k_{i}\rangle|/k_{i}\}$.
@@ VI.2 Reproducing a network spectral properties
indicating that, in such a case, the matrix $\mathbf{P}$ can be obtained as the direct product of the vector of degrees with itself. Employing the bra-ket formalism allows the calculations to be carried out quite easily, i.e. as
@@
Let us, now, consider the eMID snapshot corresponding to the day 20102010-0707-1919 and calibrate the model above on it - a choice allowing us to deal with a graphical degree sequence. Afterwards, let us generate $M=10^{3}$ configurations and treat each of them as a plausible, empirical one; solving our models on the latter leads to three distributions of $M=10^{3}$ estimates of the spectral radius each (hereby indicated with $\lambda_{1}$, to distinguish them from the reference value $\pi_{1}$ - see also Appendix E): as the left panel of fig. 3 shows, while the UBCM overestimates the spectral radius of the generative model ($\lambda_{1}^{\text{UBCM}}\simeq 9.80>\pi_{1}\simeq 9.35$) and the dcGM underestimates it ($\lambda_{1}^{\text{dcGM}}\simeq 8.66<\pi_{1}\simeq 9.35$), the fit2SM correctly reproduces it ($\lambda_{1}^{\text{fit2SM}}\simeq 9.40\gtrsim\pi_{1}\simeq 9.35$).
Figure 6: Left panel: solid lines represent the $z$-scores reading $z=(\mu-\lambda_{1}^{\text{fit2SM}})/\sigma$, while dashed lines represent the $z$-scores reading $z=(\mu-\lambda_{1})/\sigma$. In both cases, $\mu$ and $\sigma$ have been calculated via the dcGM. The similarity of the two series of values indicate that the fit2SM represents a valid generative model (e.g. for early-warning signals detection) in case the true network topology is unavailable. The pre-crisis period 20052005-$2008$ and the crisis period 20082008-$2009$ are highlighted by shaded areas. Different colours indicate different time-scale: blue-daily, magenta-weekly, green-monthly, orange-quarterly, red-yearly. Trends have been smoothed via a rolling average over the points [t10,t+10][t-10,t+10] for the daily aggregation, over the points [t7,t+7][t-7,t+7] for the weekly aggregation, over the points [t5,t+5][t-5,t+5] for the monthly aggregation, over the points [t3,t+3][t-3,t+3] for the quarterly aggregation. Right panel: irrespectively from the aggregation level at which eMID is considered, a relationship like $S=aL^{b}$ seems to hold true, the average parameters reading $a\simeq 0.51$ and $b\simeq 1.58$. Each point is a network at a given level of aggregation.
@@
  • at the daily time-scale, the fit2SM achieves the smallest average error (i.e. $0.79\pm 0.51$), performing better than the UBCM and the dcGM in $94.00\%$ and $97.60\%$ of the snapshots, respectively;
  • at the weekly time-scale, the fit2SM achieves the smallest average error (i.e. $0.97\pm 0.66$), performing better than the UBCM and the dcGM in $82.40\%$ and $98.70\%$ of the snapshots, respectively;
  • at the monthly time-scale, the fit2SM achieves the smallest average error (i.e. $0.71\pm 0.66$), performing better than the UBCM and the dcGM in $64.80\%$ and $74.50\%$ of the snapshots, respectively;
  • at the quarterly time-scale, the UBCM achieves the smallest average error (i.e. $0.43\pm 0.21$), the fit2SM performing better than the dcGM in $67.30\%$ of the snapshots but performing better than the UBCM only in $32.70\%$ of the snapshots;
  • at the yearly time-scale, the UBCM achieves the smallest average error (i.e. $0.61\pm 0.57$), the fit2SM performing better than the dcGM in $100.00\%$ of the snapshots but performing better than the UBCM only in $14.30\%$ of the snapshots.
A related comparison is depicted in the right panel of fig. 3, illustrating the distribution - summarised by the violin plot - of the relative error $(\lambda_{1}^{\text{model}}-\lambda_{1}^{\text{obs}})/\lambda_{1}^{\text{obs}}$ in reproducing the empirical spectral radius, across snapshots and temporal aggregations. It is worth noticing that the occasional, superior performance of the UBCM on the densest snapshots comes at the cost of taking the entire degree sequence as input; the fit2SM, on the other hand, solely relies upon two, global parameters, a feature making such a model substantially less demanding, in terms of informational requirements and numerical complexity.

### v4

# Reproducing the first and second moments of empirical degree distributions

@@ Abstract
The study of probabilistic models for the analysis of complex networks represents a flourishing research field. Among the former, Exponential Random Graphs (ERGs) have gained increasing attention over the years. So far, only linear ERGs have been extensively employed to gain insight into the structural organization of real-world complex networks. None, however, is capable of accounting for the variance of the empirical degree distribution. To this aim, non-linear ERGs must be considered. After showing that the usual mean-field approximation forces the degree-corrected version of the two-star model to degenerate, we define a fitness-induced variant of it. Such a ‘softened’ model is capable of reproducing the sample variance, while retaining the explanatory power of its linear counterpart, within a purely canonical framework.
@@ I Introduction
A class of models that has gained increasing attention over the years is that of Exponential Random Graphs (ERGs) [10, 16, 17, 11, 12, 18, 13, 14, 15]. ERGs belong to the category of canonical approaches, being induced by constraints that are soft, i.e., constraints that can be violated by individual configurations even if their ensemble average matches the enforced value exactly. A key advantage of canonical approaches is that the expected value of topological properties can be often expressed analytically in terms of the constraints, thereby avoiding the computational cost of generating randomized networks [19]. Microcanonical approaches, instead, artificially generate many randomized variants of the observed network, enforcing constraints that are hard, i.e., constraints that are met exactly by each graph in the ensemble [20, 21, 22, 23, 24]; this strong requirement, however, comes at the price of non-ergodicity, high computational demand, and poor generalisability [25].
So far, only linear ERGs have been extensively employed to gain insight into the structural organization of real-world systems: for simple graphs, the most important one is the Undirected Binary Configuration Model (UBCM), inducing an ensemble of configurations specified by the degree sequence; a useful fitness-induced variant of it, to be employed in presence of partial information is, instead, the density-corrected Gravity Model (dcGM), solely enforcing (a proxy of) the link density while relying on node-specific quantities identified with the node strengths $\{s_{i}\}_{i=1}^{N}$, defined as $s_{i}=\sum_{j(\neq i)}w_{ij}$, $\forall\>i$ and representing the total volume of interactions involving each node. This choice is motivated by the evidence that in economic and financial applications a node strength is observable from balance-sheets, transactional or input-output records, while the number of its counterparties (i.e. its degree) is typically not disclosed [26, 27, 28, 29, 30, 31]: treating strengths as exogenous fitnesses, thus, allows one to account for the heterogeneity of nodes without imposing any local constraint [19]: in formulas,
@@
The aim of this work is to investigate whether it is possible to define a minimal model capable of accurately reproducing both the first and second moments of empirical degree distributions, without resorting to microcanonical constraints. To this end, we consider the class of ERGs with non-linear constraints and introduce a fitness-induced variant of the two-star model that can be efficiently and reliably implemented. We test it on the transaction-level data constituting the overnight segment of the Electronic Market for Interbank Deposits (eMID), a screen-based market for unsecured deposits [8, 37, 38]. Although the raw records are directed and weighted (lender, borrower, notional amount), here, we symmetrize and binarize exposures within each, considered time window, i.e. ‘daily’, ‘weekly’, ‘monthly’, ‘quarterly’ and ‘yearly’ (the description of each aggregation procedure is reported in Appendix A).
@@ II The simplest non-linear constraint
The simplest non-linear constraint is represented by the total number of two-stars [10, 39], i.e.,
@@
Figure 1: Analysis of eMID during the 31st week (left panel) and the 42nd week (right panel) of the year 2004. Graphical representation of the agreement between the sample variance $\text{Var}[\bm{k}]$ (black, dashed, vertical line) and its expected value $\langle\text{Var}[\bm{k}]\rangle$ under the UBCM (red), the dcGM (yellow), and the fit2SM (brown): while the UBCM steadily overestimates it and the dcGM either overestimates or underestimates it, the fit2SM correctly reproduces it. Each ensemble distribution is well approximated by a Gaussian whose parameters match the corresponding average and standard deviation.
@@ III A microcanonical approach?
Since the result above suggests the microcanonical road as the only viable one, a natural choice would be that of considering the microcanonical version of the two-star model (m2SM), constraining both $L$ and $S$ exactly: since $\langle L\rangle=L$, $\text{Var}[L]=0$, and $\langle S\rangle=S$, one would obtain $\langle\text{Var}[\bm{k}]\rangle=\text{Var}[\bm{k}]$. This model is, however, homogeneous and, therefore, unable to reproduce the degree sequence.
@@
satisfying both (sets of) constraints microcanonically is equivalent to requiring that $\text{Var}[k_{i}]=0$, $\forall\>i$. In words, our results indicate that the microcanonical version of the Configuration Model (mCM) is the (simplest) one guaranteeing $\langle k_{i}\rangle=k_{i}$, $\forall\>i$ and $\langle S\rangle=S$. Moreover, as the relationship
@@
Although the (microcanonical) requirements are clear, the way to realize them is much less so; moreover, the problems affecting the microcanonical approaches mentioned in the introductory paragraph lead us to discard this route. Is a canonical way out viable?
@@ IV The mean-field approximation
the dc2SM is, however, unable to match both (sets of) constraints, as reproducing the empirical number of two-stars implies underestimating the degree of at least one node (see also Appendix C). The explanation of such a behavior lies in a simple observation: as the mean-field approximation scheme requires that
@@
letting eq. III vanish requires the generic probability coefficient of a canonical model to degenerate, becoming either 0 or 1. In other words, constraining both the degrees and the number of two-stars within the mean-field approximation scheme forces the model to become deterministic, individuating the observed configuration [^1] The asterisk indicates (the value of the quantities measured on) the observed configuration $\mathbf{A}^{*}$. $\mathbf{A}^{*}$ as the only admissible one.
@@ V A canonical way out
Let us now ask ourselves if an alternative, canonical road exists. To this aim, let us, first, notice that $\text{Var}[L]$ is divided by $N^{2}$ in eq. II: such an addendum may, thus, be expected not to play a relevant role. As a consequence, we are allowed to consider the expression:
@@
Figure 3: Left panel: distributions of the spectral radius induced by $M=10^{3}$ configurations sampled from the Chung-Lu model defined by the topology of the eMID snapshot corresponding to the day 20102010-0707-1919, according to the UBCM (red), the dcGM (yellow), and the fit2SM (brown). The three models were solved for each of the $M=10^{3}$, generated configurations and the corresponding ensembles explicitly sampled in order to obtain an estimation of $\pi_{1}$: what we find is that $\lambda_{1}^{\text{dcGM}}<\pi_{1}\lesssim\lambda_{1}^{\text{fit2SM}}<\lambda_{1}^{\text{UBCM}}$. In words, while the UBCM overestimates the spectral radius of the generative model (black, dashed, vertical line) and the dcGM underestimates it, the fit2SM correctly reproduces it. Each ensemble distribution is well approximated by a Gaussian whose parameters match the corresponding average and standard deviation. Right panel: violin plots (summarizing the distributions) of the relative error in reproducing the empirical spectral radius $\lambda_{1}^{\text{obs}}$, across snapshots and temporal aggregations. The fit2SM (brown) yields smaller errors than i) the dcGM (yellow) at all time-scales and ii) the UBCM (red) at the daily and weekly time-scales.
Constructing a fitness-induced variant of the 2SM (fit2SM) is, nevertheless, feasible: to this aim, it is enough to replace $x_{i}$ with $\sqrt{z}s_{i}$ [^3] To ease numerical manipulations, we have rescaled the strengths, dividing them by their arithmetic mean. and $k_{i}$ with its expectation under the same model in eq. 10. These substitutions lead to the generic probability coefficient reading
@@
where $\langle k_{i}\rangle_{\text{fit2SM}}\equiv\kappa_{i}$. The two, global parameters $z$ and $y$ can be estimated by solving the (system constituted by the) two non-linear coupled equations reading
@@
Figure 4: For each snapshot and aggregation level, we show the violin plots (summarising the distributions) of the relative error $\delta\text{BIC}_{\text{model}}$, defined in eq. 27. Upon remembering that a null value of such an index identifies the best-performing model on the considered snapshot, let us notice that the score induced by the fit2SM (brown) concentrates near zero - hence such a model dominates over the competing ones - at finer aggregations, while the UBCM (red) is the winner at coarser aggregations.
@@ VI Testing the performance
VI.1 Reproducing a network’s local properties
Since the degree sequence is a standard diagnostic in network reconstruction exercises, we start our analysis by testing the accuracy of the fit2SM in reproducing it. We do so by means of the quantile-quantile (QQ) plot: given the empirical sequence $\{k_{i}\}_{i=1}^{N}$ and a model-based, expected one $\{\langle k_{i}\rangle\}_{i=1}^{N}$, we sort both in increasing order and scatter the corresponding pairs of values [^4] Recall that the quantile $q$ of a set of (ordered) values indicates the percentage $q$ of smallest values. Given the empirical degree sequence and the model-based, expected one, we, thus, compute their quantile functions by sorting both in increasing order and scattering the $r$-th order statistics $\langle k\rangle_{r}$ versus $k_{r}$, $r=1,\dots,N$.; while perfect agreement would place all points on the identity line, mis-estimations of the values would be revealed by systematic deviations.
Let us, now, select five snapshots by drawing one calendar year at random and, conditionally on it, drawing one quarter, one month, one ISO week, and one trading day at random as well. Figure 2 reports the resulting QQ plots for year 2001, quarter Q3, month April, ISO week 2, and day 20012001-1111-2121; the visual inspection is accompanied by the accuracy indicators named average relative error, defined as $\text{ARE}=N^{-1}\sum_{i}|k_{i}-\langle k_{i}\rangle|/k_{i}$, and maximum relative error, defined as $\text{MRE}=\max_{i}\{|k_{i}-\langle k_{i}\rangle|/k_{i}\}$.
@@ VI.2 Reproducing a network spectral properties
indicating that, in such a case, the matrix $\mathbf{P}$ can be obtained as the direct product of the vector of degrees with itself. Employing the bra-ket formalism allows the calculations to be carried out quite easily, i.e., as
@@
Let us now consider the eMID snapshot corresponding to the day 20102010-0707-1919 and calibrate the model above on it - a choice allowing us to deal with a graphical degree sequence. Afterwards, let us generate $M=10^{3}$ configurations and treat each of them as a plausible, empirical one; solving our models on the latter leads to three distributions of $M=10^{3}$ estimates of the spectral radius each (hereby indicated with $\lambda_{1}$, to distinguish them from the reference value $\pi_{1}$; see also Appendix E): as the left panel of fig. 3 shows, while the UBCM overestimates the spectral radius of the generative model ($\lambda_{1}^{\text{UBCM}}\simeq 9.80>\pi_{1}\simeq 9.35$) and the dcGM underestimates it ($\lambda_{1}^{\text{dcGM}}\simeq 8.66<\pi_{1}\simeq 9.35$), the fit2SM correctly reproduces it ($\lambda_{1}^{\text{fit2SM}}\simeq 9.40\gtrsim\pi_{1}\simeq 9.35$).
Figure 6: Left panel: solid lines represent the $z$-scores reading $z=(\mu-\lambda_{1}^{\text{fit2SM}})/\sigma$, while dashed lines represent the $z$-scores reading $z=(\mu-\lambda_{1})/\sigma$. In both cases, $\mu$ and $\sigma$ have been calculated via the dcGM. The similarity of the two series of values indicate that the fit2SM represents a valid generative model (e.g., for early-warning signals detection) in case the true network topology is unavailable. The pre-crisis period 20052005-$2008$ and the crisis period 20082008-$2009$ are highlighted by shaded areas. Different colours indicate different time-scale: blue-daily, magenta-weekly, green-monthly, orange-quarterly, red-yearly. Trends have been smoothed via a rolling average over the points [t10,t+10][t-10,t+10] for the daily aggregation, over the points [t7,t+7][t-7,t+7] for the weekly aggregation, over the points [t5,t+5][t-5,t+5] for the monthly aggregation, over the points [t3,t+3][t-3,t+3] for the quarterly aggregation. Right panel: irrespectively from the aggregation level at which eMID is considered, a relationship like $S=aL^{b}$ seems to hold true, the average parameters reading $a\simeq 0.51$ and $b\simeq 1.58$. Each point is a network at a given level of aggregation.
@@
  • at the daily time-scale, the fit2SM achieves the smallest average error (i.e., $0.79\pm 0.51$), performing better than the UBCM and the dcGM in $94.00\%$ and $97.60\%$ of the snapshots, respectively;
  • at the weekly time-scale, the fit2SM achieves the smallest average error (i.e., $0.97\pm 0.66$), performing better than the UBCM and the dcGM in $82.40\%$ and $98.70\%$ of the snapshots, respectively;
  • at the monthly time-scale, the fit2SM achieves the smallest average error (i.e., $0.71\pm 0.66$), performing better than the UBCM and the dcGM in $64.80\%$ and $74.50\%$ of the snapshots, respectively;
  • at the quarterly time-scale, the UBCM achieves the smallest average error (i.e., $0.43\pm 0.21$), the fit2SM performing better than the dcGM in $67.30\%$ of the snapshots but performing better than the UBCM only in $32.70\%$ of the snapshots;
  • at the yearly time-scale, the UBCM achieves the smallest average error (i.e., $0.61\pm 0.57$), the fit2SM performing better than the dcGM in $100.00\%$ of the snapshots but performing better than the UBCM only in $14.30\%$ of the snapshots.
A related comparison is depicted in the right panel of fig. 3, illustrating the distribution - summarized by the violin plot - of the relative error $(\lambda_{1}^{\text{model}}-\lambda_{1}^{\text{obs}})/\lambda_{1}^{\text{obs}}$ in reproducing the empirical spectral radius, across snapshots and temporal aggregations. It is worth noticing that the occasional, superior performance of the UBCM on the densest snapshots comes at the cost of taking the entire degree sequence as input; the fit2SM, on the other hand, solely relies upon two, global parameters, a feature making such a model substantially less demanding, in terms of informational requirements and numerical complexity.